OpenStudy (anonymous):

If we let x=tan theta, then definite integral on interval [1,rt3] of sqrt(1+x^2) is equivalent to...

OpenStudy (anonymous):

If we let\[x=\tan \theta\] then \[\int\limits_{1}^{\sqrt{2}}\sqrt{1+x ^{2}}dx\] is equivalent to...

jimthompson5910 (jim_thompson5910):

sin^2(x) + cos^2(x) = 1 turns into tan^2(x) + 1 = sec^2(x) after you divide everything by cos^2(x)

jimthompson5910 (jim_thompson5910):

tan^2(x) + 1 = sec^2(x) sec^2(x) = tan^2(x) + 1 sec(x) = sqrt( tan^2(x) + 1 ) , where sec(x) > 0

OpenStudy (anonymous):

so then it's just integral of secx?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you'll have to change the limits of integration

OpenStudy (anonymous):

but how do i adjust the limits if tanx=rt3 has 2 solutions?

OpenStudy (anonymous):

i know that the bottom limit is adjusted to pi/4

jimthompson5910 (jim_thompson5910):

since sec(x) > 0, this means cos(x) > 0 hmm what to do with tan(x)

OpenStudy (anonymous):

oh wait in my answer choices only pi/3 is shown

jimthompson5910 (jim_thompson5910):

well if tan(theta) = sqrt(3), and we just look at the principal value, then theta = arctan(sqrt(3)) = pi/3

OpenStudy (anonymous):

so it has to be pi/3

jimthompson5910 (jim_thompson5910):

good, so you're going from pi/4 to pi/3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

np